### v1

# Maximally Extendable Product Codes are Good Coboundary Expanders

Abstract
We investigate the coboundary expansion property of product codes called product expansion, which plays an important role in the recent constructions of good quantum LDPC codes and classical locally testable codes. Prior research revealed that this property is equivalent to agreement testability and robust testability for products of two codes of linear distance. However, for products of more than two codes, product expansion is a strictly stronger property. In this paper, we prove that the collection of random codes over a sufficiently large field has good product expansion. We believe that in the case of four codes, these ideas can be used to construct good quantum locally testable codes in a way similar to the current constructions using only products of two codes.
Keywords— high-dimentional expanders, tensor product codes, quantum LTC
1 Introduction
The study of tensor product codes and their coboundary expansion properties has gained significant attention in recent years due to their key role in the discovery of good quantum low-density parity-check (qLDPC) codes [37] and classical locally testable codes (LTCs) [5, 37]. Already back in 2014, it was noticed that simplicial complexes with good coboundary expansion over $\mathbb{F}_{2}$, usually referred to as high-dimensional expanders (HDXs), could potentially be used to construct LTCs with large soundness [24] and qLDPC codes with large distances [23, 10].
Inspired by these works and the breakthrough paper on fiber bundle codes [20], we proposed in [37] a construction called expander lifted product code, which combines the ideas of high-dimensional expansion with the lifted product codes [36, 38] to simultaneously produce (asymptotically) good qLDPC codes and also good LTCs with large soundness.
Lifted product (LP) codes, initially introduced in QEC’2019 [36] as a natural generalization [^1] In [36] we refer to them as generalized hypergraph product codes. of hypergraph product codes [43] and other constructions [32, 18, 17, 25], were conceptualized and reformulated in [38] as a general framework to construct qLDPC codes using tensor products $\mathcal{A}\otimes_{R}\mathcal{B}$ of chain complexes over a ring $R$ such as a group algebra [^2] In fact, this idea works with arbitrary noncommutative rings $R$ (see the footnote in [38, p. 3]). $\mathbb{F}_{2}[G]$. We view vectors over $R$ as $|G|$ times larger vectors over $\mathbb{F}_{2}$ by representing each code symbol $\sum_{g\in G}a_{g}g\in R$ by the block $(a_{g})_{g\in G}$ of $|G|$ bits.
[…]
The main idea of expander LP codes [37] is to combine two (generalized) Zémor codes in a way that ensures that the corresponding 3-term chain complex has good high-dimensional expansion in both directions. Then, adopting the methods from the theory of HDXs [24, 23, 10], it is possible to show that this implies good distance of the qLDPC code and simultaneously good soundness of the LTCs from this 3-term complex. However, for this construction to work, it is required that the pair of local codes of the Zémor codes has a very specific property called product expansion that can be expressed topologically as a coboundary expansion of a complete bipartite graph naturally associated with the product code [22]. The proof in [38] shows that this local expansion lifts to the global one, which gives the desired properties. A similar approach to lift small tensor product codes to larger ones using HDXs was used in the independent work on good LTCs [5] (see also [19]). Later it was shown that expander LP codes can also be converted into other examples of good qLDPC codes [27, 6] that have more natural geometrical interpretations, and all these codes have linear time decoding algorithms [15, 28, 6, 16] as it was originally conjectured in [37].
The expansion properties of product codes in more than two dimensions are very important in the context of quantum LTCs [1]. One possible way to attack the qLTC conjecture [9], [4], [7], positing the existence of good locally testable qLDPC codes (qLTCs), is to consider a $4$-dimensional analog of the expander LP codes. One natural way is to use the tensor product of four generalized Zémor codes instead of two. In the recent result [7], it is shown that using this idea one can get almost good qLTCs provided that the collection of the four local codes and their duals both have good product expansion. There are also two more recent works relying on product expansion of more than two codes [12, 35].
In this paper, we investigate the product expansion property of several random codes. In the case of two codes, product expansion is equivalent to agreement testability and robust testability for products of two codes of linear distance. However, it is important to note that in the case of more than two codes, it is a strictly stronger property than both the agreement and the robust testability [21]. The fact that product expansion of two random codes is good was already proved in [6, 22]. Our main result in this work shows that a collection of random codes over a sufficiently large field is product-expanding with high probability. Specifically, we prove that for any fixed rates, there exists a positive constant $\rho$ such that a randomly chosen collection of codes over a sufficiently large field is $\rho$-product-expanding.
Our work builds on several new key insights and techniques. First, we establish that locally testable codes (LTCs) can be used to construct product-expanding collections of codes. We show that if a collection of codes consists of LTCs, then the product code formed by these codes is also product-expanding. However, since the duals of LTCs have small minimum distance, we cannot directly use LTCs to construct product-expanding collections of both the codes and their duals. To overcome this limitation, we introduce the concept of maximally extendable product codes, which are product codes that inherit the extendability properties of all other product codes with the same parameters. We provide a reformulation of the product expansion property in terms of extendable subsets of the dual product code, which simplifies the analysis of product expansion in higher dimensions. We prove that random codes over sufficiently large fields are maximally extendable with high probability, allowing us to construct product-expanding collections of both the codes and their duals.
In addition to our main result, we prove several useful auxiliary results. First, we show that LTCs of arbitrary length and rate exist, which allows us to perform rate adaptation in our constructions. We also prove that if we take subspaces of a product-expanding collection of codes, the resulting collection of subcodes remains product-expanding, albeit with a worse constant.
2 Definitions and main results
In this section, we define a property of a collection of codes called product expansion. Before we proceed, let us first briefly recall some standard definitions that can also be found in [22]. If $v\in\mathbb{F}_{2}^{B}$ and $A\subseteq B$ then by $v|_{A}$ we denote the restriction of $v$ to $A$. Also denote by $I_{n}$ the identity $n\times n$ matrix.
Given linear codes $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subseteq\mathbb{F}_{q}^{n}$ we can define the (tensor) product code
$\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}:=\{c\in\mathbb{F}_{q}^{[n]^{D}}\mid\forall i\in[D]\ \forall\ell\in\mathcal{L}_{i}\colon c|_{\ell}\in\mathcal{C}_{i}\},$
where $\mathbb{F}_{q}^{[n]^{D}}$ is the set of functions $c\colon[n]^{D}\to\mathbb{F}_{q}$ and $\mathcal{L}_{i}=\mathcal{L}_{i}(n,D)$ is the set of lines parallel to the $i$-th axis in the $D$-dimensional grid $[n]^{D}$, i.e.,
$\mathcal{L}_{i}:=\left\{A_{1}\times\dots\times A_{D}\subseteq[n]^{D}\mid A_{i}=[n],\left\lvert A_{j}\right\rvert=1,\forall j\neq i\right\}.$
It is convenient to introduce a notation for the dual code of a product code [44, 3]. For linear codes $\mathcal{C}_{1},\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{n}$ we denote by $\mathcal{C}_{1}\boxplus\mathcal{C}_{2}$ the code $(\mathcal{C}_{1}^{\bot}\otimes\mathcal{C}_{2}^{\bot})^{\bot}=\mathcal{C}_{1}\otimes\mathbb{F}_{q}^{n}+\mathbb{F}_{q}^{n}\otimes\mathcal{C}_{2}\subseteq\mathbb{F}_{q}^{n\times n}$. Given a collection $\mathcal{C}=(\mathcal{C}_{i})_{i\in[D]}$ of linear codes over $\mathbb{F}_{q}$, we can also define the codes
$\mathcal{C}^{(i)}:=\mathbb{F}_{q}^{n}\otimes\dots\otimes\mathcal{C}_{i}\otimes\dots\otimes\mathbb{F}_{q}^{n}=\{c\in\mathbb{F}_{q}^{[n]^{D}}\mid\forall\ell\in\mathcal{L}_{i}\colon c|_{\ell}\in\mathcal{C}_{i}\}.$
It is clear that $\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{m}=\mathcal{C}^{(1)}\cap\dots\cap\mathcal{C}^{(D)}$ and $\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}=\mathcal{C}^{(1)}+\dots+\mathcal{C}^{(D)}$. Note that every code $\mathcal{C}^{(i)}$ is the direct sum of $|\mathcal{L}_{i}|=n^{D-1}$ copies of the code $\mathcal{C}_{i}$. For $x\in\mathbb{F}_{q}^{[n]^{D}}$ we denote by $|x|_{i}$ and $\left\lVert x\right\rVert_{i}$ respectively the number and the fraction of the lines $\ell\in\mathcal{L}_{i}$ such that $a|_{\ell}\neq 0$. It is clear that $\left\lVert x\right\rVert_{i}=\frac{1}{|\mathcal{L}_{i}|}|x|_{i}$. Let us also recall that by $|x|$ and $\left\lVert x\right\rVert$ we denote respectively the Hamming weight (i.e., the number of non-zero entries) and the normalized Hamming weight (i.e., the fraction of non-zero entries) of $x$. Now we are ready to give our main definition.
Definition 2.1 (product expansion).
Given a collection $\mathcal{C}=(\mathcal{C}_{i})_{i\in[D]}$ of linear codes $\mathcal{C}_{i}\subseteq\mathbb{F}_{q}^{n}$, we say that $\mathcal{C}$ is $\rho$-product-expanding if every codeword $c\in\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}$ can be represented as a sum $c=\sum_{i\in[D]}a_{i}$ where $a_{i}\in\mathcal{C}^{(i)}$ for all $i\in[D]$, and the following inequality holds:
$\rho\sum_{i\in[D]}\left\lVert a_{i}\right\rVert_{i}\leqslant\left\lVert c\right\rVert.$ (1)
It is not hard to check that (1) can be also expressed as
$\rho n\sum_{i\in[D]}\left\lvert a_{i}\right\rvert_{i}\leqslant\left\lvert c\right\rvert.$
We can also define the product expansion factor $\rho(\mathcal{C})$ for the collection $\mathcal{C}$ as the maximal value of $\rho$ such that $\mathcal{C}$ is $\rho$-product-expanding. In [22] (see, also [6]), it is shown that a random pair of codes is $\rho$-product expanding for some positive $\rho$ depending only on the rates of these codes. In [22, Appendix B] it is also shown that the product expansion corresponds to the coboundary expansion [30, 14] in the tensor product complex obtained from the collection of codes $\mathcal{C}=(\mathcal{C}_{i})_{i\in[D]}$. This tensor product complex can be best understood as a local system (also known as a cellular sheaf) defined on the $(D-1)$-dimentional clique complex [^4] The clique complex $\mathbf{X}(\mathcal{G})$ of a graph $\mathcal{G}$ is the simplicial complex with the set of vertices $V(\mathcal{G})$ where $S\in X(\mathcal{G})$ iff the set $S$ gives a clique in $\mathcal{G}$, i.e., the vertices from $S$ are pairwise connected in $\mathcal{G}$. $\mathbf{X}(K_{n,\dots,n})$ of the complete $D$-partite graph $K_{n,\dots,n}$. With this interpretation in mind, the product expansion factor $\rho(\mathcal{C})$ can be seen as a natural generalization of the normalized Cheeger constant of a graph also known as the conductance. Indeed, in the special case when $m=2$ and the codes $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ are the repetition $[n,1,n]$ code, $\rho(\mathcal{C}_{1},\mathcal{C}_{2})$ is exactly the conductance of the bipartite graph $K_{n,n}$.
There is a degenerate case when $\mathcal{C}_{i}=\mathbb{F}_{q}^{n_{i}}$ for some $i\in[D]$ in a collection $\mathcal{C}=(\mathcal{C}_{i})_{i\in[D]}$, in which case we call this collection degenerate. In what follows we will consider only non-degenerate collections of codes.
In [22, Lemma 11] it was shown that if a non-degenerate collection $\mathcal{C}=(\mathcal{C}_{i})_{i\in[D]}$ of codes $\mathcal{C}_{i}\subsetneq\mathbb{F}_{q}^{n_{i}}$ is $\rho$-product-expanding, then each subcollection $\mathcal{C}_{I}=(\mathcal{C}_{i})_{i\in I}$, $I\subseteq[D]$, is also $\rho$-product-expanding.
Denote by ${\sf Gr}(n,k)$ the Grassmannian, i.e., the set of all linear subspaces in $\mathbb{F}_{q}^{n}$ of dimension $k$. In the next section, we will prove the following theorem.
{restatable}
theoremThMain For every tuple $(R_{1},\dots,R_{D})\in(0,1)^{D}$ there exists $\rho>0$ such that for each $n\in\mathbb{N}$ a tuple of codes $(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$ picked uniformly at random from ${\sf Gr}_{2^{t}}(n,k_{1})\times\cdots\times{\sf Gr}_{2^{t}}(n,k_{D})$, where $k_{i}\leqslant nR_{i}$, $i\in[D]$, is $\rho$-product-expanding with probability at least $1-n^{D}2^{n^{D}-t+1}$ which tends to $1$ as $t\to\infty$. Using union bound for a collection of codes and the collection of their dual codes, we obtain the following result [^5] In the special case $D=1$, we use here the fact that MDS codes with the required parameters exists for all $2^{t}>n$.
Corollary 2.2.
For each collection of intervals $I_{1},\dots,I_{D}\subseteq(0,1)$, there exists $\rho>0$ such that for all sufficiently large $n\in\mathbb{N}$ there exist codes $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subseteq\mathbb{F}_{2^{t_{n}}}^{n}$, where $t_{n}=(n+1)^{D}$, such that $\frac{1}{n}\dim\mathcal{C}_{i}\in I_{i}$ and
$\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\rho,\qquad\rho(\mathcal{C}_{1}^{\bot},\dots,\mathcal{C}_{D}^{\bot})\geqslant\rho.$
3 Proof of the main result
3.1 Locally testable codes
There are several ways how one can formally define LTCs [11]. Here, we use a rather strong form of local testability (e.g., see [26]). We say that a linear code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is $(\Delta,s)$-locally testable if it has a parity-check matrix $H\in\mathbb{F}_{q}^{m\times n}$ with rows of weight at most $\Delta$ such that for any vector $x\in\mathbb{F}_{q}^{n}$ we have
$\frac{1}{m}|Hx|\geqslant\frac{s}{n}d(x,\mathcal{C}),$
where $d(x,\mathcal{C}):=\min_{c\in\mathcal{C}}d(x,c)$, and we denote by $d(\cdot,\cdot)$ and ||$\cdot$|| ⋅ | the Hamming distance and the Hamming weight. The parameters $\Delta$ and $s$ are positive real numbers called the locality and soundness, respectively.
The next theorem shows that one can construct good LTCs of arbitrary length out of the good LTCs proposed by Dinur et al. [5]. We use the fact that the sequence of lengths in this family of good LTCs grows approximately as a geometric progression, and show that any given code length can be obtained by duplicating the codewords and adding zeros at the end.
Theorem 3.1.
For every $R\in(0,1)$ there exist constants $s>0$, $\Delta>0$, $\delta>0$ such that for all $n\in\mathbb{N}$ there is a $(\Delta,s)$-locally testable $[n,k,d]$ code for some $k\geqslant Rn$, $d\geqslant\delta n$ defined by a $\Delta$-limited parity-check matrix with $m$ rows, where [^6] This additional technical condition is used later to simplify the proof of the main result. $m\in(n/2,n]$.
3.2 Proof outline of Theorem 2.1
It is known [22] that the expansion of $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$ is almost the same as its robust testability, which in turn is much easier to prove for the product of a good code $\mathcal{C}_{1}$ and an expander LDPC code $\mathcal{C}_{2}$ (originally this result was formulated in terms of smooth codes [2]). But when we need to prove this property for both $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$ and $\mathcal{C}_{1}^{\bot}\otimes\mathcal{C}_{2}^{\bot}$, as it required in all currently existing constructions of good qLDPC codes, we cannot use LDPC codes because their dual codes have small minimum distance and, hence, $\mathcal{C}_{1}^{\bot}\otimes\mathcal{C}_{2}^{\bot}$ does not have good expansion. To build such pairs of codes, the most natural approach is to prove that pairs of random codes are product expanding, and this approach was used in [37, 6, 22]. The existing proofs for pairs of random codes are sufficiently more complicated than the proof for the case where one of the codes is a smooth code.
For $D>2$, proving the good product expansion of random codes is even more challenging. The ideas used in the proof for $D=2$ cannot be generalized in some natural way for the case $D>2$. And even for a collection of expander LDPC codes, the proof for $D=2$ cannot be directly generalized. However, the idea of using codes with some ‘‘smooth’’ properties can be applied to the collection of LTC codes. The important property of LTC codes that indirectly helps here is that if $\mathcal{C}_{1},\dots,\mathcal{C}_{D}$ are LTCs, then the code $\mathcal{C}:=\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}$ is also LTC. This property allows us to estimate the distance from a vector to the code $\mathcal{C}$ by the syndrome weight. It also helps to prove that the collection of good LTCs is product expanding (Lemmas 3.6, 3.7). Also some rate adaptation is needed to obtain examples of product expanding collection of codes with arbitrary rates (Lemma 3.8). But the duals of LTC codes are codes with small distances, so the tuple of dual codes to these LTCs cannot be product expanding. To address this, we reformulate good product expansion in terms of extendable sets in the product of dual codes (Proposition 1 and Lemmas 3.11, 3.12, 3.13) and introduce the universal property of a collection of codes that in some sense inherits all good properties from all other collections of codes [^7] It is similar to the idea of maximally recoverable (MR) codes (Definition 3.10). If a collection of codes has this universal property, then it inherits the product expansion from a collection of LTCs with the same parameters (Lemma 3.14). Then, using the Schwartz-Zippel Lemma, we prove that a random collection of codes over a large enough field has this universal property with high probability (Lemma 3.15). Hence, if the field is large enough, with high probability a collection of codes has good product expansion.
However, to implement this plan, we need to have LTCs with arbitrary length $n$ and dimension $k$. Existing constructions of LTCs [5, 37, 27, 29] can give LTCs only for some specific values of $n$ and $k$, so we need to do some rate adaptation. To do this, we first proved that we can construct good LTCs of arbitrary length [^8] It is still unknown how to construct LTCs with arbitrary dimension. and rate bounded from below (Theorem 3.1). The second step is to prove that if we reduce the dimension of the codes in a collection by taking subspaces, the product expansion does not become much worse (Lemma 3.2). In this way, we adjust the dimension of the codes without changing the length, which allows us to obtain Lemma 3.8 that states the collections of product expanding codes exist for arbitrary length and arbitrary dimensions of the codes.
We believe that LDPC codes derived from lossless expanders could replace LTCs in our proof. However, proving product expansion for such codes is more challenging than for LTCs. While this alternative strategy avoids rate adaptation and may yield better constants, it remains an open direction for future work.
3.3 Operations preserving product expansion
Lemma 3.2.
Consider codes $C_{1},\dots,C_{D}$ over $\mathbb{F}_{q}$, and a subcode $C_{1}^{\prime}\subseteq C_{1}$. Then
$\rho(C_{1}^{\prime},C_{2},\dots,C_{D})\geqslant\frac{\rho(C_{1},C_{2},\dots,C_{D})}{1+\rho(C_{2},\dots,C_{D})^{-1}}.$
Proof.
Let $\rho:=\rho(C_{1},\dots,C_{D})$, $\rho_{r}:=\rho(C_{2},\dots,C_{D})$. Consider $C_{r}:=C_{2}\boxplus\cdots\boxplus C_{D}$. Let $n_{1}$ be the length of the code $C_{1}$, $n_{r}$ be the length of the code $C_{r}$, and $n:=n_{1}n_{r}$. Consider an arbitrary codeword $w\in C_{1}^{\prime}\boxplus C_{r}$. Since $w\in C_{1}\boxplus C_{r}$, there exist words $x_{i}\in C^{(i)}$, such that $\rho\sum|x_{i}|_{i}n_{i}\leqslant|x|$.
Let $A\subseteq\mathbb{F}_{q}^{n_{1}}$ be such that $A\cap C_{1}^{\prime}=\{0\}$, $A+C_{1}^{\prime}=C_{1}$. Then $x_{1}=x_{1}^{\prime}+x_{A}$ for some $x_{1}^{\prime}\in C_{1}^{\prime}\otimes\mathbb{F}_{q}^{n_{r}},x_{A}\in A\otimes\mathbb{F}_{q}^{n_{r}}$. Since $w,x_{j},x^{\prime}_{1}\in C^{\prime}_{1}\boxplus C_{r}$ for $2\leqslant j\leqslant D$, we have $x_{1}=w-\sum_{j=2}^{D}x_{j}\in C^{\prime}_{1}\boxplus C_{r}$ and $x_{A}=x_{1}-x^{\prime}_{1}\in C^{\prime}_{1}\boxplus C_{r}$, which means $x_{A}\in(A\otimes\mathbb{F}_{q}^{n_{r}})\cap(C_{1}^{\prime}\boxplus C_{r})$. By [22, Lemma 6] we have
$(A\otimes\mathbb{F}_{q}^{n_{r}})\cap(C_{1}^{\prime}\boxplus C_{r})=(A\cap C_{1}^{\prime})\otimes\mathbb{F}_{q}^{n_{r}}+A\otimes(\mathbb{F}_{q}^{n_{r}}\cap C_{r})=A\otimes C_{r}\subseteq\mathbb{F}_{q}^{n_{1}}\otimes C_{r}.$
Therefore, $x_{A}=x_{2}^{\prime}+\cdots+x_{D}^{\prime}$ for some $x_{2}^{\prime}\in C^{(2)},\dots,x_{D}^{\prime}\in C^{(D)}$ such that $\rho_{r}\sum_{i=2}^{D}|x_{i}^{\prime}|_{i}n_{i}\leqslant|x_{A}|\leqslant|x_{A}|_{1}n_{1}\leqslant|x_{1}|_{1}n_{1}$. Put $c_{1}:=x_{1}^{\prime}$, $c_{i}:=x_{i}+x_{i}^{\prime}$ for $i=2,\dots,D$. Then $c_{1}\in C^{\prime(1)}$, $c_{i}\in C^{(i)}$ for $i=2,\dots,D$ and $w=\sum_{i=1}^{D}c_{i}$. Furthermore,
$\sum_{i=1}^{D}|c_{i}|_{i}n_{i}$ $\leqslant\sum_{i=2}^{D}|x_{i}^{\prime}|_{i}n_{i}+|x_{1}^{\prime}|_{1}n_{1}+\sum_{i=2}^{D}|x_{i}|_{i}n_{i}$
$\leqslant\frac{1}{\rho_{r}}|x_{1}|n_{1}+\sum_{i=1}^{D}|x_{i}|_{i}n_{i}\leqslant\frac{1}{\rho_{r}\rho}|x|+\frac{1}{\rho}|x|=\frac{1+\frac{1}{\rho_{r}}}{\rho}|x|.$
Hence, $\rho(C_{1}^{\prime},C_{2},\dots,C_{D})\geqslant\frac{\rho}{1+\frac{1}{\rho_{r}}}$, which completes the proof. ∎
By [22, Lemma 11] we have $\rho(C_{2},\dots,C_{D})\geqslant\rho(C_{1},\dots,C_{D})$ for codes of rates below $1$, thus we obtain the following corollary.
Corollary 3.3.
Let $C_{1},\dots,C_{D}$ be linear codes with rates below 1, $C_{i}^{\prime}\subseteq C_{i}$ for $i\in[D]$. Then
$\rho(C_{1}^{\prime},C_{2}^{\prime},\dots,C_{D}^{\prime})\geqslant 2^{-D}\bigl{(}\rho(C_{1},C_{2},\dots,C_{D})\bigr{)}^{2^{D}}.$
3.4 Product of LTCs is expanding
We say that $(\alpha_{l},\alpha_{h})$ is a soundness range of a matrix $H\in\mathbb{F}_{q}^{m\times n}$ if for any $x\in\mathbb{F}_{q}^{n}$, we have
$\alpha_{l}d(x,\ker H)\leqslant|Hx|\leqslant\alpha_{h}d(x,\ker H).$
We are interested in the case when $H$ is a parity-check matrix of an LTC, and hence $\alpha_{h}/\alpha_{l}=\Theta(1)$.
Lemma 3.4.
Let $(\alpha_{l},\alpha_{h})$ be a soundness range of a matrix $H\in\mathbb{F}_{q}^{m\times n}$, with $d(\ker H)\geq\delta n$. Then for any $A\subseteq[m]$, there exists a set $B\subseteq[n]$ such that $|B|\leq c|A|$, where $c=\frac{6}{\delta\alpha_{l}}$, and for any $y\in\operatorname{im}H$ with $\operatorname{supp}y\subseteq A$, there exists $x\in\mathbb{F}_{q}^{n}$ such that $\operatorname{supp}x\subseteq B$ and $y=Hx$.
Proof.
Fix some subset $A\subseteq[m]$, and let $a=|A|$. If $a\geq n/c$, then it suffices to choose $B=[n]$, and the lemma is proved. Thus, we can assume that $a<n/c$. Consider a basis $y_{1},\dots,y_{s}$ of the vector space $\{y\in\operatorname{im}H\ \mid\operatorname{supp}y\subseteq A\}$. For each $i\in[s]$, choose a vector $x_{i}\in\mathbb{F}_{q}^{n}$ of minimal weight such that $Hx_{i}=y_{i}$.
Since $(\alpha_{l},\alpha_{h})$ is a soundness range of $H$, we have $\left\lvert x_{i}\right\rvert\leqslant a/\alpha_{l}$, $i\in[s]$. Let $B:=\bigcup_{i=1}^{s}\operatorname{supp}x_{i}$. To prove the lemma it is enough to show that $|B|\leq 2a/\alpha_{l}<ac$. Assume the converse $|B|>2a/\alpha_{l}$, and find the smallest $s^{\prime}$ such that $|\bigcup_{i=1}^{s^{\prime}}\operatorname{supp}x_{i}|>2a/\alpha_{l}$. Consider $B^{\prime}:=\bigcup_{i=1}^{s^{\prime}}\operatorname{supp}x_{i}$. It is easy to see that $s^{\prime}>2$ and $2a/\alpha_{l}<|B^{\prime}|\leq 3a/\alpha_{l}$. Let us choose uniformly at random a vector from the linear span $\left\langle x_{1},\dots,x_{s^{\prime}}\right\rangle$. It is easy to see that the expected value of its Hamming weight is $|B^{\prime}|(1-1/q)\geq|B^{\prime}|/2>a/\alpha_{l}$, so there exists a vector $x\in\left\langle x_{1},\dots,x_{s^{\prime}}\right\rangle$ such that
$a/\alpha_{l}<|x|\leq|B^{\prime}|\leq 3a/\alpha_{l}<\frac{3n}{c\alpha_{l}}\leq\delta n/2\leq d(\ker H)/2.$
Therefore, $d(x,\ker H)=|x|$, and then $|Hx|\geq\alpha_{l}|x|>a$, which contradicts $\operatorname{supp}Hx\subseteq A$. Thus, $|B|\leq 2a/\alpha_{l}<ac$, and the lemma is proved. ∎
We will use the following obvious property of the product expansion.
Lemma 3.5.
For codes $\mathcal{C}_{1},...,\mathcal{C}_{D}\subseteq\mathbb{F}_{q}^{n}$ and zero code $\mathbf{0}\subseteq\mathbb{F}_{q}^{n}$ we have
$\rho(\mathbf{0},\mathcal{C}_{1},...,\mathcal{C}_{D})=\rho(\mathcal{C}_{1},...,\mathcal{C}_{D}).$
Lemma 3.6.
Consider codes $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subseteq\mathbb{F}_{q}^{n}$, where every code $\mathcal{C}_{i}$ has the minimum distance at least $\delta n$ and a parity-check matrix $H_{i}$ with soundness range $(\alpha_{l},\alpha_{h})$. Then for all $n$ we have that $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$ is bounded from below by a positive function $f(D,\alpha_{l},\alpha_{h},\delta)$, which does not depend on $n$.
Proof.
We will prove the lemma by induction on the dimension $D$. For $D=1$, we have $\rho(\mathcal{C}_{1})=d(\mathcal{C}_{1})/n\geqslant\delta$, so we set $f(1,\alpha_{l},\alpha_{h},\delta)=\delta$. Now, we let $D\geqslant 2$ and assume that the lemma holds for $D-1$.
Let $\mathcal{C}:=\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}$. Consider an arbitrary codeword $x\in\mathcal{C}$ and its syndrome $s:=(I_{n}\otimes H_{2}\otimes\dots\otimes H_{D})x$ for the $(D-1)$-dimensional code $\mathbf{0}\boxplus\mathcal{C}_{2}\boxplus\dots\boxplus\mathcal{C}_{D}$, where $I_{n}$ is the $n\times n$ identity matrix. Then $|s|\leqslant\alpha_{h}^{D-1}|x|$. On the other hand, $s\in\mathcal{C}_{1}\otimes\mathbb{F}_{q}^{m_{2}\times\dots\times m_{D}}$ where $m_{i}$ is the height of matrix $H_{i}$ for $i\in[D]$. Hence, $\operatorname{supp}s$ is covered by no more than $|s|/d(\mathcal{C}_{1})\leqslant|s|/(\delta n)$ lines in the first direction. Thus $\operatorname{supp}s\subseteq[n]\times A$, where $A\subseteq[m_{2}]\times\dots\times[m_{D}]$ and $|A|\leqslant|s|/(\delta n)$.
Applying Lemma 3.4 to the set $A$ and, sequentially, to the matrices $I_{n}\otimes\dots\otimes H_{i}\otimes\dots\otimes I_{n}$, $2\leqslant i\leqslant D$, we find that there exists a set $B\subseteq[n]^{D-1}$ such that $|B|\leqslant c^{D-1}|A|$, where $c=\frac{6}{\delta\alpha_{l}}$; and for any $y\in\operatorname{im}H_{2}\otimes\dots\otimes H_{D}$ with $\operatorname{supp}y\subseteq A$, there exists $z\in(\mathbb{F}_{q}^{n})^{\otimes D-1}$ such that $\operatorname{supp}z\subseteq B$ and $(H_{2}\otimes\dots\otimes H_{D})z=y$.
Consider an information set $I$ for the code $\mathcal{C}_{1}$. Denote by $s_{i}$ the slice $s|_{\{i\}\times[m_{2}]\times\dots\times[m_{D}]}$. Since $s_{i}\in\operatorname{im}H_{2}\otimes\dots\otimes H_{D}$ and $\operatorname{supp}s_{i}\subseteq A$, for all $i\in I$ there exists $z_{i}\in(\mathbb{F}_{q}^{n})^{\otimes D-1}$ such that $\operatorname{supp}z_{i}\subseteq B$ and $(H_{2}\otimes\dots\otimes H_{D})z_{i}=s_{i}$. We also define $z_{i}$ for $i\in[n]\setminus I$ by applying the systematic encoding of the code $\mathcal{C}_{1}$ for its information set $I$. It is easy to see that for $i\in[n]\setminus I$ we have $\operatorname{supp}z_{i}\subseteq B$ and $(H_{2}\otimes\dots\otimes H_{D})z_{i}=s_{i}$. Let $a^{(1)}\in(\mathbb{F}_{q}^{n})^{\otimes D}$ consist of slices $a^{(1)}|_{\{i\}\times[m_{2}]\times\dots\times[m_{D}]}=z_{i}$; then, by definition of $z_{i}$, we have $a^{(1)}\in\mathcal{C}^{(1)}$ and $|a|_{1}\leqslant|B|$.
Note that the vector $x^{\prime}:=x-a^{(1)}$ lies in the code $\mathbf{0}\boxplus\mathcal{C}_{2}\boxplus\dots\boxplus\mathcal{C}_{D}$, since
$(I_{n}\otimes H_{2}\otimes\dots\otimes H_{D})x^{\prime}=(I_{n}\otimes H_{2}\otimes\dots\otimes H_{D})x-(I_{n}\otimes H_{2}\otimes\dots\otimes H_{D})a=s-s=0.$
By Lemma 3.5 and the inductive hypothesis we have
$\rho(\mathbf{0},\mathcal{C}_{2},\dots,\mathcal{C}_{D})=\rho(\mathcal{C}_{2},\dots,\mathcal{C}_{D})\geqslant f(D-1,\alpha_{l},\alpha_{h},\delta).$
Hence $x^{\prime}=\sum_{j=2}^{D}a^{(j)}$ for some $a^{(j)}\in\mathcal{C}^{(j)}$, $2\leqslant j\leqslant D$ such that
$nf(D-1,\alpha_{l},\alpha_{h},\delta)\sum_{j=2}^{D}|a^{(j)}|_{j}\leqslant|x^{\prime}|.$
We have:
$|a^{(1)}|\leqslant|B|n\leqslant c^{D-1}|A|n\leqslant c^{D-1}|s|/\delta\leqslant c^{D-1}\alpha_{h}^{D-1}\delta^{-1}|x|.$
$|x^{\prime}|=|x-a^{(1)}|\leqslant|x|+|a^{(1)}|\leqslant\left(1+c^{D-1}\alpha_{h}^{D-1}\delta^{-1}\right)|x|$
Therefore,
$x=a^{(1)}+x^{\prime}=\sum_{j=1}^{n}a^{(j)},\quad a^{(j)}\in\mathcal{C}^{(j)}\mbox{ for }j\in[D].$
We can estimate $N:=\sum_{j=1}^{D}|a^{(j)}|_{j}$ as follows:
$N\leqslant|B|+\frac{|x^{\prime}|}{nf(D-1,\alpha_{l},\alpha_{h},\delta)}\leqslant\frac{(c\alpha_{h})^{D-1}}{\delta n}|x|+\left(1+\frac{(c\alpha_{h})^{D-1}}{\delta}\right)\frac{|x|}{nf(D-1,\alpha_{l},\alpha_{h},\delta)}.$
Using that $c\geqslant 1,\delta\leqslant 1,\alpha_{h}\geqslant 1,f(D-1,\alpha_{l},\alpha_{h},\delta)\leqslant\delta\leqslant 1$, simplify and substitute $c$ from Lemma 3.4:
$N\leqslant\frac{|x|}{n}\cdot\frac{3(c\alpha_{h})^{D-1}}{\delta\cdot f(D-1,\alpha_{l},\alpha_{h},\delta)}=\frac{3\left(\frac{6\alpha_{h}}{\alpha_{l}\delta}\right)^{D-1}}{\delta\cdot f(D-1,\alpha_{l},\alpha_{h},\delta)}\cdot\frac{|x|}{n}.$
Thus, we can take $f(D,\alpha_{l},\alpha_{h},\delta):=\frac{\delta\cdot f(D-1,\alpha_{l},\alpha_{h},\delta)}{3\left(\frac{6\alpha_{h}}{\alpha_{l}\delta}\right)^{D-1}}$, and the lemma is proved. ∎
Note that in the proof we did not use soundness range for the code $\mathcal{C}_{1}$, therefore $\mathcal{C}_{1}$ can be arbitrary code with the distance at least $\delta n$.
The following statement directly follows from the definition of LTCs, the $\Delta$-limited matrix, and the soundness range.
Lemma 3.7.
If a code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is $(\Delta,s)$-locally testable with a $\Delta$-limited parity-check matrix $H$ of size $m\times n$, where $n/2\leqslant m\leqslant n$, then $(s/2,\Delta)$ is a soundness range of the matrix $H$.
Lemma 3.8.
For arbitrary $r_{1},\dots,r_{D}\in(0,1)$, there exists $\rho>0$ such that for any $n\in\mathbb{N}$ and $k_{i}\in\mathbb{N}$ such that $k_{i}\leqslant r_{i}n$, there exist codes $\mathcal{C}_{i}\subseteq\mathbb{F}_{2}^{n}$, $i\in[D]$, such that $\dim\mathcal{C}_{i}=k_{i}$ and
$\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\rho.$
Proof.
Let $r:=\max(r_{1},\dots,r_{D})$. By Theorem 3.1, for the rate $r$, choose $s>0,\Delta>0,\delta>0$. Let $\rho:=2^{-D}(f(D,s/2,\Delta,\delta))^{2^{D}}$, where the function $f$ is from Lemma 3.6. Fix $n$ and $k_{i}$ such that $k_{i}\leqslant r_{i}n$. The required collection of codes will be constructed in 4 steps.
  1. 1.
    By Theorem 3.1, there exists a $(\Delta,s)$-LTC code $\mathcal{C}$ of length $n$ such that $\dim\mathcal{C}\geqslant rn$, $d(C)\geqslant\delta n$, with a $\Delta$-limited parity-check matrix $H$ of size $m\times n$, where $n/2\leqslant m\leqslant n$.
  2. 2.
    By Lemma 3.7, the pair $(s/2,\Delta)$ is a soundness range of the matrix $H$.
  3. 3.
    By Lemma 3.6, with $\alpha_{l}=s/2,\alpha_{h}=\Delta$, we have
    $\rho(\underbrace{\mathcal{C},\dots,\mathcal{C}}_{D\text{ times}})\geqslant f(D,s/2,\Delta,\delta)>0.$
  4. 4.
    For each code $i\in[D]$, we have $\dim\mathcal{C}\geqslant rn\geqslant k_{i}$; choose an arbitrary subcode $\mathcal{C}_{i}\subseteq\mathcal{C}$ of dimension $k_{i}$. By Corollary 3.3, we have
    $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant 2^{-D}\rho(\mathcal{C},\dots,\mathcal{C})\geqslant 2^{-D}(f(D,s/2,\Delta,\delta))^{2^{D}}=\rho.$
The lemma is proved. ∎
3.5 Maximally Extendable Product Codes
Given the grid $[n]^{D}$, we consider the set of axis-parallel lines $\mathcal{L}(n,D):=\bigcup_{i=1}^{D}\mathcal{L}_{i}(n,D)$, where $\mathcal{L}_{i}(n,D)$ are the lines in the $i$-th direction. If we have a code $\mathcal{C}=\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}$, $\mathcal{C}_{i}\subseteq\mathbb{F}_{q}^{n}$, then to a line $\ell\in\mathcal{L}_{i}(n,D)$ we can assign the code
$\mathcal{C}_{\ell}:=\{c\in\mathbb{F}_{q}^{[n]^{D}}\mid\operatorname{supp}c\subseteq\ell,c|_{\ell}\in\mathcal{C}_{i}\}.$
The $\varepsilon$-closure of the set $M\subseteq[n]^{D}$ is defined as the minimum set $[M]_{\varepsilon}\subseteq[n]^{D}$ such that for any line $\ell\in\mathcal{L}(n,D)$, either $\ell\subseteq[M]_{\varepsilon}$ or $|\ell\cap[M]_{\varepsilon}|<\varepsilon n$. If $[M]_{\varepsilon}=M$, then the set $M$ is called $\varepsilon$-closed.
Let $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subseteq\mathbb{F}_{q}^{n}$, $M\subseteq[n]^{D}$. Consider the set $L(M):=\left\{\ell\in\mathcal{L}(n,D)\mid\ell\subseteq M\right\}$ of all axis-parallel lines contained in $M$.
Definition 3.9 (inner-generated sets).
We call a set $M$ inner-generated for a code $\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$ if for each its codeword $c$ with $\operatorname{supp}c\subseteq M$ we have $c\in\sum_{\ell\in L(M)}{\mathcal{C}_{\ell}}$, i.e., it can be represented as a sum of codewords along lines contained in $M$.
Since $\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$ is dual for the code $\mathcal{C}^{\mathrm{T}}:=\mathcal{C}_{1}^{\perp}\otimes\ldots\otimes\mathcal{C}_{D}^{\perp}$ we can view its codewords as checks for $\mathcal{C}^{\mathrm{T}}$. With this interpretation in mind, the set of bit positions $M\subseteq[n]^{D}$ is inner-generated iff every check $z$ of $\mathcal{C}^{\mathrm{T}}$ with $\operatorname{supp}z\subseteq M$ can be obtained as a linear combination of the local checks from $\mathcal{C}_{\ell}$, $\ell\subseteq M$. The last condition is in turn equivalent to the condition that every local codeword $c_{M}\in\mathbb{F}_{q}^{M}$ satisfying all local checks $z|_{M}$, where $z\in\mathcal{C}_{\ell}$, $\ell\subseteq M$, can be extended to a global codeword $c\in\mathcal{C}^{\mathrm{T}}$, $c|_{M}=c_{M}$, satisfying all the checks $z\in\mathcal{C}^{\mathrm{T}}$. We say that a set $M\subseteq[n]^{D}$ with this property is extendable in the product code $\mathcal{C}^{\mathrm{T}}$. Thus we have the following fact.
Proposition 1.
A set $M\subseteq[n]^{D}$ is inner-generated for $\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$ iff $M$ is extendable in $\mathcal{C}_{1}^{\perp}\otimes\ldots\otimes\mathcal{C}_{D}^{\perp}$.
Since a collection of codes $(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$ has good product expansion when every codeword from $\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$ can be represented as a sum of a sufficiently small number of codewords from $\mathcal{C}_{\ell}$, $\ell\in\mathcal{L}(n,D)$, then, taking into account Proposition 1, it is very beneficial to have as many extendable sets $M$ as possible in the product code $\mathcal{C}_{1}^{\perp}\otimes\ldots\otimes\mathcal{C}_{D}^{\perp}$. This leads us to the following definition.
Definition 3.10 (maximally extendable product code).
We say that a product code $\mathcal{C}=\otimes_{i\in[D]}\mathcal{C}_{i}\subseteq\mathbb{F}_{2^{t}}^{[n]^{D}}$ is maximally extendable if for every other product code $\mathcal{C}^{\prime}=\otimes_{i\in[D]}\mathcal{C}^{\prime}_{i}\subseteq\mathbb{F}_{2^{t^{\prime}}}^{[n]^{D}}$ with $\dim\mathcal{C}_{i}=\dim\mathcal{C}_{i}^{\prime}$, $i\in[D]$, when $M$ is extendable in $\mathcal{C}^{\prime}$ it is also extendable in $\mathcal{C}$.
As we will prove later, every product code $\mathcal{C}=\otimes_{i\in[D]}\mathcal{C}_{i}$ has good expansion provided that the code $\mathcal{C}^{\mathrm{T}}=\otimes_{i\in[D]}\mathcal{C}_{i}^{\perp}$ is maximally extendable. The following lemma shows that $\rho$-product expansion implies that all $\rho$-closed sets are inner-generated.
Lemma 3.11.
Let $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subsetneq\mathbb{F}_{q}^{n}$, $\rho:=\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$. Then each $\rho$-closed subset $M\subseteq[n]^{D}$ is inner-generated for the code $\mathcal{C}:=\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$.
Proof.
For a set $M\subseteq[n]^{D}$ denote $\mathcal{C}_{L(M)}:=\sum_{\ell\in L(M)}\mathcal{C}_{\ell}$. Suppose there exists $\rho$-closed set $M\subseteq[n]^{D}$ which is not inner-generated. Then the set $S:=(\mathcal{C}\cap\mathbb{F}_{q}^{M})\setminus\mathcal{C}_{L(M)}$ is not empty. Consider a vector $x\in S$ of minimal weight. Since $\rho=\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$, the vector $x$ can be represented as a sum $\sum_{i=1}^{s}x_{i}$ of $s$ words $x_{i}\in\mathcal{C}_{\ell_{i}}$ along some lines $\ell_{i}\in\mathcal{L}(n,D)$ where $s\leqslant|x|/(\rho n)$. Without loss of generality, we can assume that the lines ordered in such a way that $\ell_{i}\in L(M)$ for $1\leqslant i\leqslant s^{\prime}$ and $\ell_{i}\not\in L(M)$ for $s^{\prime}<i\leqslant s$. Let $x^{\prime}:=\sum_{i=1}^{s^{\prime}}x_{i}$, $x^{\prime\prime}:=x-x^{\prime}=\sum_{i=s^{\prime}+1}^{s}x_{i}$. Then $\operatorname{supp}x^{\prime}\subseteq M$, therefore $\operatorname{supp}x^{\prime\prime}\subseteq M$. Since $x\not\in\mathcal{C}_{L(M)}$, $x^{\prime}\in\mathcal{C}_{L(M)}$ we have $x^{\prime\prime}\not\in\mathcal{C}_{L(M)}$. Hence $x^{\prime\prime}\in S$. Since the set $M$ is $\rho$-closed and $\ell_{i}\not\in L(M)$ for $s^{\prime}<i\leqslant s$, we have $|\ell_{i}\cap M|<\rho n$. Taking into account that $\operatorname{supp}x^{\prime\prime}\subseteq M$, we obtain
$|x^{\prime\prime}|\leqslant\sum_{i=s^{\prime}+1}^{s}\bigl{|}x_{i}|_{M}\bigr{|}\leqslant\sum_{i=s^{\prime}+1}^{s}|\ell_{i}\cap M|<(s-s^{\prime})\rho n\leqslant s\rho n\leqslant|x|.$
Thus, we have a contradiction with the assumption that $x$ has minimal weight in the set $S$. This completes the proof. ∎
Lemma 3.12.
For any $\varepsilon>0$ and $D\in\mathbb{N}$, there exists a constant $c>0$ such that $|[M]_{\varepsilon}|\leqslant c|M|$ for any $n\in\mathbb{N}$ and $M\subseteq[n]^{D}$.
Proof.
Consider a set $M$. We are going to construct a set $M^{\prime}$ that includes $[M]_{\varepsilon}$. Set $\varepsilon^{\prime}:=\varepsilon/2^{D}$. For a subset $I\subseteq[D]$, let $M_{I}$ denote the set of cells covered by $|I|$-dimensional hyperplanes that are parallel to the axes corresponding to the indices in $I$, containing at least $(\varepsilon^{\prime}n)^{|I|}$ elements of $M$. Then $M^{\prime}:=\bigsqcup_{I\subseteq[D]}M_{I}$ contains $M$.
Let us show that $[M]_{\varepsilon}\subseteq M^{\prime}$. For this, it is enough to verify that the set $M^{\prime}$ is $\varepsilon$-closed. Consider an arbitrary line $\ell\in\mathcal{L}(n,D)$ such that $\ell\not\subseteq M^{\prime}$. Suppose that $|\ell\cap M_{I}|\geqslant\varepsilon^{\prime}n$ for some $I\subseteq[D]$. Then in the $(|I|+1)$-dimensional hyperplane $P$, passing through the line $\ell$ in directions from the set $I$, there are at least $\varepsilon^{\prime}n$ $|I|$-dimensional hyperplanes from $M_{I}$, meaning $|P\cap M|\geqslant\varepsilon^{\prime}n\cdot(\varepsilon^{\prime}n)^{|I|}=(\varepsilon^{\prime}n)^{|I|+1}$. Therefore, $P\in M_{I\cup\{j\}}$, where $j$ is the direction of line $\ell$, meaning $\ell\subseteq P\subseteq M^{\prime}$, a contradiction. Thus, $|\ell\cap M_{I}|<\varepsilon^{\prime}n$. Then $|\ell\cap M^{\prime}|\leqslant\sum_{I\subseteq[D]}|\ell\cap M_{I}|<2^{D}\varepsilon^{\prime}n=\varepsilon n$. Since the line $\ell$ was arbitrary and not entirely contained in $M^{\prime}$, the set $M^{\prime}$ is $\varepsilon$-closed.
Now, it remains to estimate $|M^{\prime}|$. It is easy to see that $|M_{I}|\leqslant(|M|/(\varepsilon^{\prime}n)^{|I|})n^{|I|}\leqslant|M|(\varepsilon^{\prime})^{-|I|}$. Therefore,
$|[M]_{\varepsilon}|\leqslant|M^{\prime}|\leqslant\sum_{I\subseteq[D]}|M|(\varepsilon^{\prime})^{-|I|}=|M|(1+1/\varepsilon^{\prime})^{D}\leqslant|M|\left(\frac{2^{D}+1}{\varepsilon}\right)^{D}.$
Setting $c=\left(\frac{2^{D}+1}{\varepsilon}\right)^{D}$, we get the statement of the lemma. ∎
Lemma 3.13.
Let $\mathcal{C}_{1},\dots,\mathcal{C}_{D}\subsetneq\mathbb{F}_{q}^{n}$, $\varepsilon>0$ and each $\varepsilon$-closed subset $M\subseteq[n]^{D}$ is inner-generated for code $\mathcal{C}:=\mathcal{C}_{1}\boxplus\ldots\boxplus\mathcal{C}_{D}$. Then $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\gamma(\varepsilon,D)$ where $\gamma(\varepsilon,D):=\frac{\varepsilon^{D}}{D(2^{D}+1)^{D}}$.
Proof.
Consider arbitrary codeword $x\in\mathcal{C}$ and let $M:=[\operatorname{supp}x]_{\varepsilon}$. By Lemma 3.12 we have $|M|\leqslant c|x|$ where $c=\left(\frac{2^{D}+1}{\varepsilon}\right)^{D}$. Since $M$ is $\varepsilon$-closed, it is inner-generated, hence $x=\sum_{\ell\in L(M)}x_{\ell}$ for some $x_{\ell}\in\mathcal{C}_{\ell}$. Therefore, $x$ is represented as a sum of at most $|L(M)|$ codewords along lines contained in $M$. Since each line has size $n$, and each point belongs to at most $D$ lines, we have
$|L(M)|\leqslant|M|\frac{D}{n}\leqslant cD\frac{|x|}{n}=\frac{1}{\gamma(\varepsilon,D)}\frac{|x|}{n}.$
Therefore, $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\gamma(\varepsilon,D)$. ∎
Lemma 3.14.
For all $D\in\mathbb{N}$ there is a function $\mu_{D}\colon(0,1)^{D}\to(0,1)$ such that for every tuple of rates $(r_{1},\dots,r_{D})\in(0,1)^{D}$ and maximally extendable code $\mathcal{C}_{1}^{\perp}\otimes\dots\otimes\mathcal{C}_{D}^{\perp}$ such that $\mathcal{C}_{i}\in{\sf Gr}_{2^{t}}(n,k_{i})$ and $k_{i}\leqslant r_{i}n$ we have $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\mu_{D}(r_{1},\dots r_{D})$.
Proof.
Let us fix some $D\in\mathbb{N}$ and tuple of rates $(r_{1},\dots,r_{D})\in(0,1)^{D}$. Let $\hat{\rho}$ be a constant from Lemma 3.8 for rates $r_{1},\dots,r_{D}$. Put $\mu_{D}(r_{1},\dots,r_{D}):=\gamma(\hat{\rho},D)$ where $\gamma$ is the function from Lemma 3.13.
Consider a maximally extendable code $\mathcal{C}_{1}^{\perp}\otimes\dots\otimes\mathcal{C}_{D}^{\perp}$ satisfying the conditions of the lemma. By Lemma 3.8 there exist codes $\hat{\mathcal{C}}_{i}\in{\sf Gr}_{2}(n,k_{i})$, $i\in[D]$ such that $\rho(\hat{\mathcal{C}}_{1},\dots,\hat{\mathcal{C}}_{D})\geqslant\hat{\rho}$. By Lemma 3.11 each $\hat{\rho}$-closed subset $M\in[n]^{D}$ is inner generated for the code $\hat{\mathcal{C}}:=\hat{\mathcal{C}}_{1}\boxplus\dots\boxplus\hat{\mathcal{C}}_{D}$. Since the product code $\mathcal{C}_{1}^{\perp}\otimes\dots\otimes\mathcal{C}_{D}^{\perp}$ is maximally extendable, by Proposition 1 each $\hat{\rho}$-closed subset $M\in[n]^{D}$ is also inner generated for the code $\mathcal{C}:=\mathcal{C}_{1}\boxplus\dots\boxplus\mathcal{C}_{D}$. Therefore, by Lemma 3.13 we have $\rho(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\geqslant\gamma(\hat{\rho},D)=\mu_{D}(r_{1},\dots,r_{D})$. ∎
The next lemma follows from Theorem 1 in [39], but for the reader’s convenience, we provide a direct proof here, which also comes with better bounds. As in [39] we adopt the approach used in [13, Lemma 32] to show the existence of maximally-recoverable codes.
Lemma 3.15.
For a collection $(\mathcal{C}_{1},\dots,\mathcal{C}_{D})$ picked uniformly at random from ${\sf Gr}_{2^{t}}(n,k_{1})\times\cdots\times{\sf Gr}_{2^{t}}(n,k_{D})$ the code $\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}$ is maximally extendable with probability at least $1-n^{D}2^{n^{D}-t+1}$.
Proof.
It is clear that the parity-check matrix $H$ for the code $\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}$ can be obtained by stacking the matrices $\otimes_{i\in[D]}M_{i}$, $j\in[D]$, where $M_{j}=H_{j}$, and $M_{i}=I_{n}$ for $i\neq j$ (each such matrix checks the constraints of the product code in the $j$-th direction).
Given a collection of free variables $\mathbf{p}$, let $\mathbb{F}_{2}[\mathbf{p}]$ and $\mathbb{F}_{2}(\mathbf{p})$ denote respectively the ring of polynomials and the field of rational functions in the variables from $\mathbf{p}$. Consider a parametrization for the codes $\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}$ by polynomial parity-check matrices $H_{i}\in\mathbb{F}_{2}[\mathbf{p}]^{(n-k_{i})\times n}$, where $(H_{i})_{u,v}:=p_{i,u,v}$, and $\mathbf{p}:=(p_{i,u,v})_{i\in[D],u\in[n-k_{i}],v\in[n]}$ is the collection of $N:=n\sum_{i=1}^{D}(n-k_{i})$ free parameters we used.
For a subset of indices $S\subseteq[n]^{D}$ we consider two submatrices $H_{S}$ and $H^{S}$ of $H$. The matrix $H_{S}$ is obtained from $H$ by retaining only the columns corresponding to the codeword symbols with indices in $S$, while $H^{S}$ is composed of the rows from $H_{S}$ corresponding to the parity-check equations in $H$ involving only such symbols. We say that a vector $\mathbf{a}\in\mathbb{F}_{2^{t}}^{N}$ is a good substitution for the polynomial matrix $H$ if for each $S\subseteq[n]^{D}$, we get:
$\operatorname{rk}H^{\prime}(\mathbf{a})=\operatorname{rk}H^{\prime};H^{\prime}\in\{H_{S},H^{S}\},$ (2)
where the left side rank is taken over the finite field $\mathbb{F}_{2^{t}}$ and the right side over the infinite field $\mathbb{F}_{2}(\mathbf{p})$. Let us denote by $U(2^{t})$ the set of all good substitutions over the field $\mathbb{F}_{2^{t}}$.
Claim 1.
If $\mathbf{a}\in U(2^{t})$, then $\mathcal{C}=\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}$ is maximally extendable, where $\mathcal{C}_{i}=\ker H_{i}(\mathbf{a})$, $i\in[D]$.
Proof.
[…]
Figure 1: $S$ is extendable in $\mathcal{C}=\ker H(\mathbf{a})$ iff $\ker H_{S}(\mathbf{a})=\mathcal{C}|_{S}$ iff any information set $J$ of $\ker H^{S}(\mathbf{a})$ is contained in some information set $I$ of $\ker H(\mathbf{a})$. In this case, the gray submatrices of $H(\mathbf{a})$ and $H^{S}(\mathbf{a})$ both must be full-rank.
The main idea of the proof is a simple observation that $S\subseteq[n]^{D}$ is extendable in $\mathcal{C}$ iff $\mathcal{C}|_{S}=\ker H^{S}(\mathbf{a})$, where $\mathcal{C}|_{S}=\{c|_{S}\mid c\in\mathcal{C}\}$ represents the vectors in $\mathbb{F}_{2^{t}}^{S}$ that can be extended to full codewords, and the local code $\ker H^{S}(\mathbf{a})$ represents the vectors in $\mathbb{F}_{2^{t}}^{S}$ satisfying all local checks imposed by the codes $\mathcal{C}_{\ell}$, $\ell\subseteq S$. Alternatively, $S$ is extendable iff every information set $J$ of the local code $\ker H^{S}(\mathbf{a})$ can be extended to an information set $I$ of the global code $\mathcal{C}=\ker H(\mathbf{a})$ (see Fig. 1).
It is easy to see that the inclusion $\mathcal{C}|_{S}\subseteq\ker H^{S}(\mathbf{a})$ always holds. Thus, in order to show that set $S$ is extendable in $\mathcal{C}$, it is enough to check that
$\dim\mathcal{C}|_{S}\geqslant\dim\ker H^{S}(\mathbf{a}).$ (3)
First, note that $\mathcal{C}=\ker H(\mathbf{a})$ is maximally recoverable. This means that for any other product code $\ker H(\mathbf{a}^{\prime})$ with the same dimension, if $I$ is an information set for $\ker H(\mathbf{a}^{\prime})$, then $I$ is also an information set for $\mathcal{C}$. This follows directly from the fact that $\mathbf{a}\in U(2^{t})$, and substituting parameters into a polynomial matrix does not increase its rank:
$\operatorname{rk}H_{[n]^{D}\setminus I}(\mathbf{a}^{\prime})\leqslant\operatorname{rk}H_{[n]^{D}\setminus I}=\operatorname{rk}H_{[n]^{D}\setminus I}(\mathbf{a})\leqslant n^{D}-|I|.$
For a good substitution $\mathbf{a}\in U(2^{t})$, the dimension of the local code $\ker H^{S}(\mathbf{a})$ is minimal among all possible substitutions $\mathbf{a}^{\prime}$, since $\operatorname{rk}H^{S}(\mathbf{a}^{\prime})\leqslant\operatorname{rk}H^{S}=\operatorname{rk}H^{S}(\mathbf{a})$. At the same time, the dimension of the projection $\mathcal{C}|_{S}$ is maximal, as it equals the largest size of a set $J\subseteq S$ that can be extended to an information set of $\mathcal{C}$. This maximality is guaranteed because $\mathcal{C}$ is maximally recoverable.
Finally, if $S$ is extendable in another product code $\mathcal{C}^{\prime}=\ker H(\mathbf{a}^{\prime})$ with $\dim\mathcal{C}^{\prime}=\dim\mathcal{C}$, then $S$ is also extendable in $\mathcal{C}$. This follows because $\mathcal{C}^{\prime}|_{S}=\ker H^{S}(\mathbf{a}^{\prime})$, and the maximality of $\dim\mathcal{C}|_{S}$ combined with the minimality of $\dim\ker H^{S}(\mathbf{a})$ implies
$\dim\mathcal{C}|_{S}\geqslant\dim\mathcal{C}^{\prime}|_{S}=\dim\ker H^{S}(\mathbf{a}^{\prime})\geqslant\dim\ker H^{S}(\mathbf{a}),$
therefore (3) holds for $S$ and $\mathcal{C}|_{S}=\ker H^{S}(\mathbf{a})$. Thus, $S$ is extendable in $\mathcal{C}$, proving that $\mathcal{C}$ is maximally extendable.
Claim 2.
$\mathsf{P}\left\{\mathbf{a}\in U(2^{t})\right\}\geqslant 1-n^{D}2^{n^{D}-t+1}.$
Proof.
Note that we have $2^{n^{D}+1}$ polynomial matrices $H_{S}$ and $H^{S}$ for all possible $S\subseteq[n]^{D}$. Let us arbitrary choose one maximal non-singular submatrix in each such matrix, and denote by $f$ the product of their determinants. From conditions (2) we see that for $\mathbf{a}\in\mathbb{F}^{N}_{2^{t}}$ we have $f(\mathbf{a})\neq 0$ iff $\mathbf{a}\in U(2^{t})$. It is also clear that $\deg f\leqslant n^{D}2^{n^{D}+1}$. Thus, by the Schwartz–Zippel lemma (e.g., see [34, Theorem 7.2]), the probability that for a uniformly random $\mathbf{a}\in\mathbb{F}^{N}_{2^{t}}$ we have $\mathbf{a}\not\in U(2^{t})$ is bounded above as $\deg f/2^{t}\leqslant n^{D}2^{n^{D}-t+1}$, and the claim is proved. ∎
To complete the prove of the lemma we also need to show that the uniform distribution over codes is approximately the same as the uniform distribution over parity-check matrices. This follows from the following facts:
  • If we pick a code $\mathcal{C}\in{\sf Gr}(n,k)$ uniformly at random, and pick a matrix $H$ for $\mathcal{C}$ uniformly at random from the set of $(n-k)\times n$ parity-check matrices of the code $\mathcal{C}$, then the distribution of matrices $H$ obtained this way is uniform over all $(n-k)\times n$ matrices of maximum rank.
  • For a good substitution $\mathbf{a}$, all matrices $\ker H_{i}(\mathbf{a})$ have maximal rank, thus
    $\mathsf{P}\left(\mathbf{a}\in U(2^{t})\mid\dim\ker H_{i}(\mathbf{a})=k_{i}\mbox{ for }i\in[D]\right)\geqslant\mathsf{P}\left(\mathbf{a}\in U(2^{t})\right)\geqslant 1-n^{D}2^{n^{D}-t+1}.$
  • The parameter vector $\mathbf{a}$ corresponds exactly to the elements of all parity-check matrices $H_{i}$.
Thus, if we pick uniformly at random a tuple of codes $(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\in{\sf Gr}_{2^{t}}(n,k_{1})\times\cdots\times{\sf Gr}_{2^{t}}(n,k_{D})$ and then pick uniformly at random parity-check matrices $H_{1},\dots,H_{D}$ for these codes, with probability at least $1-n^{D}2^{n^{D}-t+1}$ the corresponding substitution $\mathbf{a}$ is good. Hence, by Claim 1, the code $\mathcal{C}_{1}\otimes\dots\otimes\mathcal{C}_{D}$ is maximally extendable. ∎
[…]
Proof.
Put $\rho:=\mu_{D}(R_{1},\dots,R_{D})$ where $\mu$ is the function from Lemma 3.14. Fix some numbers $n\in\mathbb{N}$, $k_{i}\leqslant R_{i}n$, $t\in\mathbb{N}$. By Lemma 3.14, if for a collection of codes $(\mathcal{C}_{1},\dots,\mathcal{C}_{D})\in{\sf Gr}_{2^{t}}(n,k_{1})\times\cdots\times{\sf Gr}_{2^{t}}(n,k_{D})$ the code $\mathcal{C}_{1}^{\perp}\otimes\dots\otimes\mathcal{C}_{D}^{\perp}$ is maximally extendable, then this collection is $\rho$-product-expanding. Moreover, by Lemma 3.15, if we pick these codes uniformly at random, the probability of this event is at least $1-n^{D}2^{n^{D}-t+1}\to 1$ as $t\to\infty$. This completes the proof. ∎
Appendix A LTCs of arbitrary length
In this section, we give the proof of Theorem A.1 by constructing good LTCs of arbitrary length out of the good LTCs proposed by Dinur et al. [5]. However, first, we give an extended formulation of Theorem 1.1 from [5], where we also include additional information about the constructed LTCs that appeared in its proof.
Theorem A.1.
For every $r\in(0,1)$ there exists constants $s>0$, $\delta>0$, $\Delta\in\mathbb{N}$ and $q\in\mathbb{N}$ such that for all $i\in\mathbb{N}$ there exists $(\Delta,s)$-locally testable $[n_{i},k_{i},d_{i}]$ codes with $\Delta$-limited parity-check matrix $H_{i}$ for some $k_{i}\geqslant rn_{i}$ and $d_{i}\geqslant\delta n_{i}$, where $n_{i}:=\frac{1}{8}\Delta^{2}(q^{3i}-q^{i})$.
Remark A.2.
In Theorem 1.1 of [5], LTCs are constructed using the square complex $\mathrm{Cay}(A,G,B)$ with two generating sets $A$ and $B$ of size $\Delta$, where the bits are assigned to squares and the checks to the edges. The row weights in the parity-check matrix $H_{i}$ of the obtained $[n_{i},k_{i},d_{i}]$ do not exceed $\Delta$. Each local code has the rate no less than $3/4$, so its parity-check matrix has no more than $\Delta/4$ rows, and each square participates in $4$ local codes corresponding to the edges of the square. Therefore, each square is involved in no more than $\Delta$ checks, thus the code matrix is $\Delta$-limited. It will be convenient for us next to choose an overcomplete parity-check matrix for local codes by duplicating some of its rows multiple times so that it has exactly $\left\lfloor\Delta/4\right\rfloor$ rows; in this case, the matrix $H_{i}$ of the entire LTC code will have the width $n$ and the height $m=\frac{4\left\lfloor\Delta/4\right\rfloor}{\Delta}n$. Assuming that the rate is $<1$, we get $\Delta\geq 4$, $n/2<m\leq n$.
Remark A.3.
It is also worth noting that the test in the article [5], which checks the entire set of rows in the parity-check matrix corresponding to the product of local codes, is slightly different from our definition of a local code, where exactly one row of the parity-check matrix is checked. Therefore, our parameters may differ by a constant factor from the parameters of the test from [5]. In particular, according to our definition, the locality of the constructed code is $\Delta$, while in the test from [5], the locality is $\Delta^{2}$. Additionally, our soundness parameter $s$ differs from the parameter $\kappa$ from [5].
We will also need a lemma that the direct sum of locally testable codes is also a locally testable code.
Lemma A.4.
Let $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ be an $(\Delta,s)$-locally testable code of length $n$ with $\Delta$-limited parity-check matrix $m\times n$. Then for any $t\in\mathbb{N}$ the code $\mathcal{C}\otimes\mathbb{F}_{q}^{t}$ is also an $(\Delta,s)$-locally testable code with $\Delta$-limited parity-check matrix $tm_{0}\times tn_{0}$.
Proof.
Let $H$ be the parity-check matrix of code $\mathcal{C}$. Consider the $m^{\prime}\times n^{\prime}$ parity-check matrix of code $\mathcal{C}^{\prime}:=\mathcal{C}\otimes\mathbb{F}_{q}^{t}$:
$H^{\prime}:=H\otimes I_{t}=\begin{pmatrix}H_{0}&0&\cdots&0\\ 0&H_{0}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&H_{0}\end{pmatrix},$
where $I_{t}$ is the identity matrix of size $t\times t$, $m^{\prime}=tm$, $n^{\prime}=tn$.
Consider an arbitrary $w=(x_{1},\ldots,x_{t})\in\mathbb{F}_{q}^{n^{\prime}}$, $x_{i}\in\mathbb{F}_{q}^{n}$ for $i\in[t]$. Then we have
$d(w,\mathcal{C}^{\prime})$ $=\sum_{i=1}^{t}d(x_{i},\mathcal{C}_{0})\leqslant\sum_{i=1}^{t}\frac{1}{s}\cdot\frac{n}{m}|Hx_{i}|=\frac{1}{s}\cdot\frac{n^{\prime}}{m^{\prime}}\sum_{i=1}^{t}|Hx_{i}|=\frac{1}{s}\cdot\frac{n^{\prime}}{m^{\prime}}|H^{\prime}w|.$
Furthermore, it is easy to see that the maximum weight of rows and columns in the matrix $H$ is the same as in the matrix $H$. The lemma is proved. ∎
Additionally, we will need a lemma that local testability is preserved when adding a zero code.
Lemma A.5.
Consider a $(\Delta,s)$-locally testable code $\mathcal{C}\subsetneq\mathbb{F}_{q}^{n}$ with a $\Delta$-limited $m\times n$ parity-check matrix $H$, $n/2\leqslant m\leqslant n$, and let $\mathbf{0}_{t}\subseteq\mathbb{F}_{q}^{t}$ be the zero code of length $t$. Then the code $\mathcal{C}^{\prime}:=\mathcal{C}\oplus\mathbf{0}_{t}$ is a $(\Delta,s^{\prime})$-locally testable code with a $\Delta$-limited parity-check matrix where $s^{\prime}:=\min(\frac{s^{\prime}}{2},1)$.
Proof.
Consider the parity-check matrix of the code $\mathcal{C}^{\prime}$:
$H^{\prime}:=\begin{pmatrix}H&0\\ 0&I_{t}\end{pmatrix},$
where $I_{t}$ is the identity matrix of size $t\times t$. Let $n^{\prime}:=n+t$, $m^{\prime}:=m+t$ be the width and height of matrix $H^{\prime}$. From the constraints on $m,n$, it is easy to see that $1\leqslant\frac{n^{\prime}}{m^{\prime}}\leqslant\frac{n}{m}\leqslant 2\frac{n^{\prime}}{m^{\prime}}$.
Consider an arbitrary vector $w=(x,y)\in\mathbb{F}_{q}^{n^{\prime}}$, $x\in\mathbb{F}_{q}^{n}$, $y\in\mathbb{F}_{q}^{t}$. Then we have:
$d(w,\mathcal{C}^{\prime})$ $=d(x,\mathcal{C})+|y|\leqslant\frac{1}{s}\cdot\frac{n}{m}|Hx|+|y|$
$=\frac{1}{s}\cdot 2\frac{n^{\prime}}{m^{\prime}}|Hx|+\frac{n^{\prime}}{m^{\prime}}|I_{t}y|$
$\leqslant\max\left(\frac{2}{s},1\right)\frac{n^{\prime}}{m^{\prime}}(|Hx|+|y|)=\frac{1}{s^{\prime}}\cdot\frac{n^{\prime}}{m^{\prime}}|H^{\prime}w|.$
It is easy to see that the weight of columns and rows of matrix $H^{\prime}$ is bounded by $\max(\Delta,1)=\Delta$. The lemma is proved. ∎
Proof of Theorem 3.1.
Let $r:=(1+R)/2$. According to Theorem A.1, there exist constants $s_{0}>0,\delta_{0}>0,\Delta\in\mathbb{N},q\in\mathbb{N}$ such that for any $i\in\mathbb{N}$ there exists a $(\Delta,s)$-locally testable $[n_{i},k_{i},d_{i}]$ code $\mathcal{C}_{i}$ with a $\Delta$-limited parity-check matrix $H_{i}$ of size $m_{i}\times n_{i}$ (according to Remark A.2, we can ensure that $n_{i}/2<m_{i}\leqslant n_{i}$), where $n_{i}=\frac{1}{8}\Delta^{2}(q^{3i}-q^{i})$, $k_{i}\geqslant rn_{i}$, $d_{i}\geqslant\delta n_{i}$.
Let us define
$\delta:=\min\left(\frac{1-r}{n_{1}},\delta_{0}\frac{1-r}{q^{3}+q}\right),$ $s:=\min\left(\frac{1-r}{n_{1}},\frac{s_{0}}{2}\right),$
and show that for any $n\in\mathbb{N}$ there exists a code with the required parameters.
Fix an arbitrary $n\in\mathbb{N}$. If $n<n_{1}/(1-r)$, then the code $\mathcal{C}=\mathbb{F}_{2}^{n}$ with rate 1 has distance $1>\delta n$. This trivial code is $(\Delta,s)$-locally testable since the distance from any vector to the code is zero.
Next, consider the main case $n\geqslant n_{1}/(1-r)$. Note that for all $i\in\mathbb{N}$ the following holds
$n_{i+1}/n_{i}=q\frac{q^{2i+2}-1}{q^{2i}-1}\leq q(q^{2}+1).$
Let us put
$j:=\max\left\{i\in\mathbb{N}\mid n_{i}\leqslant(1-r)n\right\}.$
Then $j\geqslant 1$ and $n_{j}\leqslant n(1-r)<n_{j+1}$, which implies $n_{j}\geqslant\frac{n_{j+1}}{q^{3}+q}>\frac{1-r}{q^{3}+q}n$. Let $n=n_{j}t+u$, $t\in\mathbb{N}$, $0\leqslant u<n_{j}$. Consider the code $\mathcal{C}^{\prime}:=\mathcal{C}_{j}\otimes\mathbb{F}_{q}^{t}$ and $\mathcal{C}:=\mathcal{C}^{\prime}\oplus\mathbf{0}_{u}$. Let us find the parameters of the code $\mathcal{C}$. It is easy to see that the code $\mathcal{C}^{\prime}$ has length $tn_{j}=n-u$, and the code $\mathcal{C}$ has length $n$. For the dimension $k:=\dim\mathcal{C}$, we have
$k=\dim\mathcal{C}^{\prime}=t\dim\mathcal{C}_{j}=tk_{j}\geqslant tn_{j}r=nr-ur>nr-n(1-r)r=nr^{2}>n(2r-1)=nR.$
For the distance we have
$d(\mathcal{C})=d(\mathcal{C}^{\prime})=d(\mathcal{C}_{j})\geqslant\delta_{0}n_{j}>\delta_{0}\frac{1-r}{q^{3}+q}n=\delta n.$
By Lemma A.4, code $\mathcal{C}^{\prime}$ is a $(\Delta,s_{0})$-locally testable code with a $\Delta$-limited parity-check matrix $H^{\prime}$ of size $m^{\prime}\times n^{\prime}$, $n^{\prime}/m^{\prime}=n_{j}/m_{j}\in[1,2)$. Therefore, by Lemma A.5, the code $\mathcal{C}$ is a $(\Delta,\min(1,\frac{s_{0}}{2}))$-locally testable code, and its matrix is also $\Delta$-limited. Thus, we have constructed a code with the required parameters, and the theorem is proved. ∎